For a certain casino slot machine, the odds in favor of a win

are given as 71 to 29. Express the indicated degree of likelihood as a probability value between 0 and 1 inclusive.

The probability is ?

​(Round to two decimal places as​ needed.)
Answer #1

the odds in favor of a win are given as 71 to 29

probability = 71 / (71+29) = 71 / 100

probability = 0.71

probability value between 0 and 1 inclusive.